OpenStudy (anonymous):

The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work

OpenStudy (imstuck):

Solving for a means that you have to get a alone. Begin by subtracting the ut over to the d side. This gives you\[d-ut=.5at ^{2}\]Now divide by the .5t^2 to get this:\[\frac{ d-ut }{ .5t ^{2}}=a\]

OpenStudy (imstuck):

And that's it!

OpenStudy (anonymous):

2(d-ut)=at2 2(d-ut)/t^2=A?;)

OpenStudy (anonymous):

So a= d-ut/.dt2?

OpenStudy (imstuck):

Looks that way to me (except you forgot your .5)

OpenStudy (anonymous):

um ya pretty much

OpenStudy (anonymous):

Okay, Thank you!

OpenStudy (anonymous):

sure:)